Using balmer series calculate ionisation energy for hydrogen aton in kj mol per second
1
Expert's answer
2017-03-10T10:33:05-0500
Since we are dealing with hydrogen, we know the equation is E=-RH(Z^2/n^2) or E=-RH(1/1). The rydberg constant is 2.178E-18, so take that number, multiply it by 6.022E23 (a mole) and divide it by 1000 (KJ) = 1312 kJ/mol.